Helium-filled balloons are used to carry scientific instruments high into the atmosphere. Suppose a balloon is launched when the temperature is 22.5 °C, and the barometric pressure is 751 mm Hg. If the balloon’s volume is 4.34×104 L (and no helium escapes from the balloon), what will the volume be at a height of 20 miles, where the pressure is 76.0 mm Hg, and the temperature is -33.0 °C
Answer #1

Solution:

According to ideal gas equation,

P1V1/T1 = P2V2/T2

Where,

P1 = 751 mm Hg

V1 = 4.34 x 10^4 L

T1 = 22.5 + 273 = 295.5

P2 = 76 mm Hg

T2 = 273 -33 = 240 K

V2 = ?

Therefore,

V2 = P1V1T2 / P2T1

= 751 x 4.34 x 10^4 x 295.5 / 76 x 240

= 5.28 x 10^5 L

Thus, volume will be 5.28 x 10^5 L